Question: Write an equation in slope - intercept form of the line that passes through (1, -1), and has a slope of 5?
Somehow when I did it I had gotten the answer Y+1=5(x-1) and it was wrong. So if you could help that would be great. Thank you

Answer by mananth(16946)

The equation of line is given by
y-y1=m(x-x1)
where (x1,y1) are the co ordinates and m the slope
(1,-1) & m=5
y-(-1)=5(x-1)
y+1=5x-5
y=5x-6